OpenStudy (anonymous):

an article sold at a certain fixed price. by selling it at 2/3 of that price, one loses 10%. the gain % on selling it at original price is ? a.20 b.33 1/3 c.35 d.40

OpenStudy (zepp):

If he sells at 2/3 of the price, that means that 1/3 is the 10% he lost.

OpenStudy (zepp):

We are interested in the gain, 1/3 would be the gain, if he sells at the original price, therefore gain = 33 1/3 %

OpenStudy (anonymous):

how u saying 1/3 is gain?

OpenStudy (zepp):

He sells at 2/3, original price is 1 2/3 -> 1 = +1/3

OpenStudy (anonymous):

if he sells at 2/3rd of the price, it leads to 10% lost.. so if he sells for original price, he wont get 1/3rd gain.. am i right??

OpenStudy (anonymous):

Think about it this way. P is price. M is margin (what you gain). (P + M) * 2/3 = P * (1-0.1) P(1-0.1) means you only get back 90% of what you paid for it. You lose 10. So you solve for M.

OpenStudy (anonymous):

Should leave you with M = 7/5 P, which is the same as M = 140/100 P, or M = 140% P Your gain should be 40%

OpenStudy (zepp):

Umm, didn't think of that D:

OpenStudy (anonymous):

I understand, its not very intuitive :(

OpenStudy (anonymous):

ans s 35%

OpenStudy (anonymous):

let cp = 100 sp * 2/3 = 90 so sp = (90*2)/3 = 135 so 35% gain... but i cant understand it properly

OpenStudy (anonymous):

Define cp, sp.

OpenStudy (anonymous):

cost price, selling price

OpenStudy (anonymous):

Its not that straight forward. The cost price is P. The selling price is P plus a margin M. The margin is your profit. So from your problem you get that (P + M), when that is 2/3, you get back 90% of P. Which means 0 margin, and 10% less of what you paid for. So 10% loss. In the end the equation should be \[(P + M ) {2 \over 3} = P * 0.9\]

OpenStudy (anonymous):

Which you want to solve for M, as above

OpenStudy (anonymous):

Let the original selling price be : x And now the new selling price is : (2/3) of x = 2x/3 Now the formula for Cost Price is: \[CP = \frac{100}{(100 - Loss)} \times SP\] Where SP is Selling Price.. Here, Loss = 10% and SP = 2x/3 Put in the formula for find the Cost Price: \[CP = \frac{100}{90} \times \frac{2x}{3}\] \[CP = \frac{20x}{27}\] This is the Cost Price for that article.. Selling Price is : x (we have assumed it earlier) Gain = Cost Price - Selling Price \[Gain = x - \frac{20x}{27}\] \[Gain = \frac{7x}{27}\] Now find the Gain Percentage: Gain% = \[= \frac{\frac{7x}{27}}{\frac{20x}{27}} \times 100 = \frac{7x \times 27}{27 \times 20x}\] = 35%

OpenStudy (anonymous):

Sorry, there I have written wrong.. Gain = Selling Price - Cost Price...